# CPSDBench: A Large Language Model Evaluation Benchmark for Chinese Public Security Domain thanks: Citation: Xin Tong, Bo Jin. CPSDBench: A Large Language Model Evaluation Benchmark and Baseline for Chinese Public Security Domain. DOI:000000

### 11 Feb 2024 v1

@@ Abstract
Keywords Large Language Model  $\cdot$ Public Security  $\cdot$ Evaluation Benchmark
@@ 3.1 Tasks and Datasets
To comprehensively evaluate the performance of LLMs in the field of public security, CPSDBench has constructed a four-dimensional assessment task system encompassing text classification, information extraction, question answering, and text generation. These tasks incorporate a variety of datasets collected from actual public security business scenarios, thereby enabling a more thorough and in-depth assessment of LLMs’ understanding and generation capabilities concerning public security-related data. Considering the high costs associated with utilizing commercial-grade LLMs (such as ChatGPT, ChatGLM-130B), the size of each dataset is typically limited to a range of several dozen to several hundred samples.
@@ 3.3 Evaluation Metrics
For text classification tasks, we have chosen Accuracy, Precision, Recall, and F1-Score as evaluation metrics. The computation of this metric is illustrated in Equations (1).
@@
For information extraction tasks, the typical metrics are Precision, Recall, and F1-Score. The computation of these metrics is illustrated in Equations (2)-(4). However, these metrics often focus on calculating results that are exact matches. The output of LLMs may differ literally from the labels but be semantically equivalent. For instance, the label for a time entity in a sample might be ’9:00 am on August 6, 2021’, while the LLM’s output could be ’around 9:00 am on August 6, 2021’, where the additional word ’around’ also originates from the input sample. From a human perspective, such LLM predictions should also be considered correct. Moreover, there are more complex scenarios in Chinese. In light of this, we have designed a hybrid evaluation metric. It comprises two steps: firstly, calculating the exact match score between the LLM’s output and the label. Secondly, for predictions that are not exact matches, we calculate the Levenshtein distance, as shown in Equation (5), and deem predictions above a certain threshold as correct for calculating a fuzzy match score. Finally, we obtain a comprehensive score by weighting these two types of scores, as depicted in Equation (6). In the equation, Whereas, a and b represent two input strings, and i and j respectively denote the indices of string a and string b. Additionally, for entities requiring high precision, such as the name and amount entities in telecommunication fraud case entity recognition, we have only calculated the exact match score, disregarding the fuzzy similar results.
@@
In the objective reading comprehension component of the question-answering task, this study adopts Accuracy as the primary evaluation metric to quantify the model’s performance in identifying the correct answers. For generative question-answering and text generation tasks, the study selects Bilingual Evaluation Understudy (BLEU) and Recall-Oriented Understudy for Gisting Evaluation (ROUGE) as assessment metrics. These metrics are utilized to evaluate the model’s performance in terms of the quality and accuracy of generated text. The specific computational processes are
defined in Equations (7) and (8), respectively. Equation (7) elaborates on the calculation method for the BLEU metric, focusing primarily on the lexical overlap between the generated text and the reference text. Equation (8) describes the calculation process for the ROUGE metric, emphasizing the extent to which the generated text covers and retains the content of the original text. Together, these two metrics provide a comprehensive quantitative basis for evaluating the performance of models in generative question-answering and text generation tasks.Furthermore, we have also selected the BERT Score, calculated based on the BERT-base model, as the evaluation criterion for subjective reading comprehension questions. This approach aims to measure the semantic similarity between LLMs’ predictions and the labels from a semantic perspective.
@@ 4.1 Performance Evaluation
In this study, we conducted a comprehensive evaluation of advanced LLMs to explore their performance across different natural language processing tasks. Through systematic assessments in four key tasks: text classification, information extraction, question answering, and text generation, our aim was to unveil the strengths and limitations of mainstream LLMs in public security tasks.
@@
Furthermore, our analysis also compared the performance disparities between proprietary and open-source models. Generally, proprietary models outperform open-source models. This can be attributed to proprietary models typically having a larger number of parameters and access to more extensive private training datasets, factors that collectively influence the final performance of the models.
Lastly, we analyzed the impact of parameter scale on LLMs. By comparing the performance of different parameter scale versions of the Atom and XVerse models, we found that the number of parameters plays a key role in enhancing natural language understanding capabilities. However, this impact is less pronounced in text generation tasks. Surprisingly, the 13B version of XVerse even outperformed its 65B version in text generation tasks, suggesting that besides model size, other factors such as model architecture optimization, training strategies, or data quality may have a significant impact on the performance of LLMs.
@@ 4.2 Error Analysis
To gain a deeper understanding of the limitations of existing LLMs in handling complex tasks, we conducted a detailed error analysis of examples that performed poorly in the CPSDBench tasks. These errors can primarily be categorized into the following four types, reflecting challenges in comprehension, generation, and safety aspects of LLMs:
@@
(1) Triggering Sensitive Alerts. We observed that LLMs tend to trigger safety mechanisms when processing tasks containing sensitive content. Taking iFLYTEK Spark as an example, its built-in stringent user input review mechanism resulted in the model nearly refusing to answer all questions in tasks such as telecommunication fraud detection. This excessive cautiousness led to extremely low accuracy rates in such tasks, for instance, only a 2.5% accuracy rate in the telecommunication fraud detection task. This indicates that LLMs might experience performance degradation in tasks involving sensitive information due to overly stringent safety filtering.
@@
(2) Incorrect Output Format. Issues with output format represent one of the common error types for LLMs. In text classification and information extraction tasks, incorrect output formats directly impact the calculation of evaluation metrics. For example, the QWEN and ERNIE models incorporated additional explanatory information in their outputs, which did not comply with the task requirements for concise output formats. Moreover, when faced with open-ended questions, models tend to generate complete sentences rather than the required keywords or brief answers, leading to discrepancies between the predicted results and the standard answers.
@@
(3) Inability to Understand Instructions. We discovered that models with smaller parameter sizes, such as Atom-1B, often struggle to accurately comprehend user instructions. These models may merely repeat the content of the prompt in text generation tasks rather than generating new, relevant information. This indicates that a model’s comprehension and generative abilities are closely related to its parameter scale, with smaller models facing limitations in understanding complex instructions or generating pertinent content.
@@
(4) Errors in Output Content. Regarding errors in output content, especially in information extraction tasks, some models exhibited issues with missing key entities. Moreover, in text classification tasks, some models like Atom-1B tended to excessively output labels of a specific category, such as classifying instances as "normal" in telecommunication fraud detection tasks. This may reflect biases in the model’s training or overfitting to the training data distribution.
@@ Acknowledgments
This was was supported in part by National Social Science Foundation Key Project of China (No.20AZD114).

### 03 Mar 2024 v2

# School of Information and Network Security

@@ Abstract
Keywords Large Language Models  $\cdot$ Evaluation Benchmark  $\cdot$ Public Security  $\cdot$ Police
@@ 3 Method
In China, public safety constitutes the core of police work, encompassing critical areas such as crime prevention, case investigation, intelligence analysis, and the maintenance of public order. This significant responsibility is typically borne by various specialized police forces to ensure targeted and efficient maximization. For instance, narcs play an indispensable role in drug control efforts, with their primary tasks being the identification of individuals, locations, and items related to drug activities, thereby advancing drug control initiatives. Cyber police focus on maintaining order in cyberspace, analyzing and assessing the public’s emotional responses to social events, timely identifying and dealing with online rumors, and ensuring the stability and safety of the online environment. criminal police, security officers, and traffic police play crucial roles in the investigation of criminal cases, handling civil disputes, and the control of traffic violations, respectively. Furthermore, in response to the severe challenges posed by the emerging crime of telecommunications fraud, the criminal police department has established an anti-fraud center. Its main tasks include quickly filtering potential fraudulent information, effectively issuing warnings, and blocking telecommunications fraud activities to protect the public’s financial safety and reduce economic losses. This collaborative division of labor ensures the comprehensiveness and professionalism of public safety work, providing a solid foundation for maintaining social stability and citizen safety.
@@ 3.1 Tasks and Datasets
To comprehensively evaluate the performance of LLMs in the field of public security, CPSDBench has constructed a four-dimensional assessment task system encompassing text classification, information extraction, question answering, and text generation. These tasks incorporate a variety of datasets collected from actual public security business scenarios, thereby enabling a more thorough and in-depth assessment of LLMs’ understanding and generation capabilities concerning public security-related data. The evaluative content of CPSDBench and its correspondence to the police category in China are illustrated in the figure. Considering the high costs associated with utilizing commercial-grade LLMs (such as ChatGPT, ChatGLM-130B), the size of each dataset is typically limited to a range of several dozen to several hundred samples.
Figure 1: The evaluative content of CPSDBench and its correspondence to the police category in China.
@@ 3.3 Evaluation Metrics
For text classification tasks, we have chosen Accuracy, Precision, Recall, and F1-Score as evaluation metrics. The computation of this metric is illustrated in Equation ( 1).
@@
For information extraction tasks, the typical metrics are Precision, Recall, and F1-Score. The computation of these metrics is illustrated in Equations ( 2)-( 4). However, these metrics often focus on calculating results that are exact matches. The output of LLMs may differ literally from the labels but be semantically equivalent. For instance, the label for a time entity in a sample might be ’9:00 am on August 6, 2021’, while the LLM’s output could be ’around 9:00 am on August 6, 2021’, where the additional word ’around’ also originates from the input sample. From a human perspective, such LLM predictions should also be considered correct. Moreover, there are more complex scenarios in Chinese. In light of this, we have designed a hybrid evaluation metric. It comprises two steps: firstly, calculating the exact match score between the LLM’s output and the label. Secondly, for predictions that are not exact matches, we calculate the Levenshtein distance, as shown in Equation ( 5), and deem predictions above a certain threshold as correct for calculating a fuzzy match score. Finally, we obtain a comprehensive score by weighting these two types of scores, as depicted in Equation ( 6). In the equation, Whereas, a and b represent two input strings, and i and j respectively denote the indices of string a and string b. Additionally, for entities requiring high precision, such as the name and amount entities in telecommunication fraud case entity recognition, we have only calculated the exact match score, disregarding the fuzzy similar results.
@@
In the objective reading comprehension component of the question-answering task, this study adopts Accuracy as the primary evaluation metric to quantify the model’s performance in identifying the correct answers. For generative question-answering and text generation tasks, the study selects Bilingual Evaluation Understudy (BLEU) and Recall-Oriented Understudy for Gisting Evaluation (ROUGE) as assessment metrics. These metrics are utilized to evaluate the model’s performance in terms of the quality and accuracy of generated text. The specific computational processes are defined in Equations ( 7) and ( 8), respectively. Equation ( 7) elaborates on the calculation method for the BLEU metric, focusing primarily on the lexical overlap between the generated text and the reference text. Equation ( 8) describes the calculation process for the ROUGE metric, emphasizing the extent to which the generated text covers and retains the content of the original text. Together, these two metrics provide a comprehensive quantitative basis for evaluating the performance of models in generative question-answering and text generation tasks.Furthermore, we have also selected the BERT Score, calculated based on the BERT-base model, as the evaluation criterion for subjective reading comprehension questions. This approach aims to measure the semantic similarity between LLMs’ predictions and the labels from a semantic perspective.
@@ 4.1 Performance Evaluation
In this study, we conducted a comprehensive evaluation of advanced LLMs to explore their performance across different natural language processing tasks. The results are shown in Table 2-Table 5.Through systematic assessments in four key tasks: text classification, information extraction, question answering, and text generation, our aim was to unveil the strengths and limitations of mainstream LLMs in public security tasks.To achieve a more intuitive and comprehensive comparison of the performance of LLMs, this study also utilized radar charts to visually present the overall efficacy of the models, as shown in Figure 3.
@@
Our analysis also compared the performance disparities between proprietary and open-source models. Generally, proprietary models outperform open-source models. This can be attributed to proprietary models typically having a larger number of parameters and access to more extensive private training datasets, factors that collectively influence the final performance of the models.
Furthermore, we analyzed the impact of parameter scale on LLMs. By comparing the performance of different parameter scale versions of the Atom and XVerse models, we found that the number of parameters plays a key role in enhancing natural language understanding capabilities. However, this impact is less pronounced in text generation tasks. Surprisingly, the 13B version of XVerse even outperformed its 65B version in text generation tasks, suggesting that besides model size, other factors such as model architecture optimization, training strategies, or data quality may have a significant impact on the performance of LLMs.
Finally, input length is an important factor affecting the performance of LLMs in public safety tasks. We observed the variation in F1 scores of LLMs when processing texts of different lengths in the task of extracting information from drug-related cases, as illustrated in Figure 1. The analysis indicates that for the majority of models, the predictive capability of LLMs tends to decline with an increase in the length of the input text.
Figure 3: Overall performance comparison of LLMs. The text classification score is obtained by calculating the average accuracy across three subtasks. For information extraction, the F1 score from drug-related case tasks is used as the final score. The text generation task employs the BERT score as its final score.
@@
Figure 4: The impact of text length on the F1 scores for LLMs. To avoid the occurrence of negative values in the violin plot due to the kernel density estimation process, a phased processing approach was adopted for visualization. Due to the poor performance of the Atom-1B model in information extraction tasks, it is not shown in the figure.
@@ 4.2 Error Analysis
To gain a deeper understanding of the limitations of existing Large Language Models (LLMs) in handling complex tasks, we conducted a comprehensive error analysis on examples that performed poorly in the CPSDBench tasks. These errors can primarily be categorized into four types: triggering sensitive alerts, incorrect output formatting, failure to understand instructions, and erroneous content generation. These categories reflect the challenges LLMs face in comprehension, generation, and safety aspects. Figure 5 and figure 6 illustrate the distribution of error types in LLMs across two representative tasks: sentiment classification and telecommunications fraud detection.
Figure 5: Error Type Distribution of LLMs in Sentiment Classification Task.
Figure 6: Error Type Distribution of LLMs in Telecommunications Fraud Detection Task.
4.2.1 Triggering Sensitive Alerts
@@
We observed that LLMs tend to trigger safety mechanisms when processing tasks containing sensitive content. Taking iFLYTEK Spark as an example, its built-in stringent user input review mechanism resulted in the model nearly refusing to answer all questions in tasks such as telecommunication fraud detection. This excessive cautiousness led to extremely low accuracy rates in such tasks, for instance, only a 2.5% accuracy rate in the telecommunication fraud detection task. This indicates that LLMs might experience performance degradation in tasks involving sensitive information due to overly stringent safety filtering.
4.2.2 Incorrect Output Format
@@
Issues with output format represent one of the common error types for LLMs. In text classification and information extraction tasks, incorrect output formats directly impact the calculation of evaluation metrics. For example, the ChatGLM and ERNIE models incorporated additional explanatory information in their outputs, which did not comply with the task requirements for concise output formats. Moreover, when faced with open-ended questions, models tend to generate complete sentences rather than the required keywords or brief answers, leading to discrepancies between the predicted results and the standard answers.To assess the impact of neglecting formatting errors on the performance of LLMs) this study implemented an accuracy calculation method based on the principle of fuzzy matching. Through this method, we observed significant variations in the accuracy of multiple LLMs in tasks such as rumor detection and telecommunications fraud identification, with specific changes illustrated in figure  7 and figure  8. Notably, some models experienced an increase in accuracy by more than 20% after adopting fuzzy matching. This demonstrates that improving the formatting output capabilities of LLMs holds significant importance for tasks related to public safety.
Figure 7: After adopting fuzzy accuracy as the evaluation metric, the performance of LLMs in rumor detection tasks changed. The accuracy of the ChatGLM-4 model increased by 28%.
Figure 8: Following the implementation of fuzzy accuracy as an assessment criterion, the effectiveness of LLMs in telecommunications fraud detection tasks altered. The accuracy of the XVERSE2-13B model improved by 21%.
4.2.3 Inability to Understand Instructions
@@
We discovered that models with smaller parameter sizes, such as Atom-1B, often struggle to accurately comprehend user instructions. These models may merely repeat the content of the prompt in text generation tasks rather than generating new, relevant information. This indicates that a model’s comprehension and generative abilities are closely related to its parameter scale, with smaller models facing limitations in understanding complex instructions or generating pertinent content.
4.2.4 Errors in Output Content
@@
Regarding errors in output content, especially in information extraction tasks, some models exhibited issues with missing key entities. Moreover, in text classification tasks, some models like Atom-1B tended to excessively output labels of a specific category, such as classifying instances as "normal" in telecommunication fraud detection tasks. This may reflect biases in the model’s training or overfitting to the training data distribution.Figures  9,  10, and  11 respectively illustrate the output bias of LLMs in three text classification tasks.
Figure 9: In the execution of sentiment classification tasks, the majority of models more frequently predict emotional tendencies as either "positive" or "negative," rather than opting for a "neutral" label.
Figure 10: Predictive bias of LLMs in rumor detection tasks
Figure 11: In telecommunications fraud detection tasks, LLMs exhibit a tendency to output "malicious" labels, which may lead to a higher false positive rate when applying them to this task.
@@ Acknowledgments
This study was supported in part by National Social Science Foundation Key Project of China (No.20AZD114).